OpenStudy (anonymous):

divide (9x^2-1)/(8x-4) by (3x^2+5x-2)/(2x^2+1x-1) to its simplest form and explain the process please! :) I'm slowly catching on but I still need some guidance

OpenStudy (anonymous):

\(\frac a b\) divided by \(\frac c d \) is \(\frac {ad}{bc} \). Now refer to Calli's answer to your earlier question.

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {a \over b} \div {x \over y} = {a \over b} \times {y \over x}}\) Notice how nothing happens to a/b but x/y becomes the reciprocal

OpenStudy (anonymous):

i need help factoring since when you multiply the two problems together after flipping one of the fractions you get 18^4+1x+1 / 3x^2+40x+8 AHHHHH